OpenStudy (anonymous):

maki made 1/2 of a pound of trail mix. if she puts 1/6 of a pound into each bag, how many bags can maki fill

OpenStudy (anonymous):

@emmathis

OpenStudy (texaschic101):

(1/2) / (1/6) = 1/2 * 6/1 = 6/2 = 3 check.. 3 * 1/6 = 1/2 3/6 = 1/2 1/2 = 1/2 (correct) 3 bags can be filled

OpenStudy (anonymous):

thanks